# A Comprehensive Survey on World Models for Embodied AI

### v1

@@ 1 Introduction
Embodied AI aims to equip agents to perceive complex, multimodal environments, act within them, and anticipate how their actions will alter future world states [1, 2]. Central to this capability is the world model, an internal simulator that captures environment dynamics to support both forward and counterfactual rollouts for perception, prediction, and decision making [3, 4]. This survey focuses on world models that yield actionable predictions for embodied agents, distinguishing them from static scene descriptors or purely generative visual models that do not capture controllable dynamics.
@@
Cognitive science suggests humans construct internal models of the world by integrating sensory inputs. These models do not merely predict and simulate future events but also shape perception and guide action [5, 6, 7]. Motivated by this view, early AI research on world models was rooted in model-based reinforcement learning (RL), where latent state-transition models were used to improve sample efficiency and planning performance [8]. The seminal work of Ha and Schmidhuber [9] crystallized the term world model and inspired the Dreamer series [10, 11, 12], highlighting how learned dynamics can drive imagination-based policy optimization. More recently, advances in large-scale generative modeling and multimodal learning have expanded world models beyond their initial focus on policy learning into general-purpose environment simulators capable of high-fidelity future prediction, exemplified by models like Sora [13] and V-JEPA 2 [14]. This expansion has diversified functional roles, temporal modeling strategies, and spatial representations, while introducing inconsistencies in terminology and taxonomy across sub-communities.
@@
Faithfully capturing environment dynamics requires addressing both the temporal evolution of states and the spatial encoding of scenes [3]. Long-horizon rollouts are susceptible to error accumulation, which establishes coherence as a central challenge in video prediction and policy imagination [15, 16]. Similarly, coarse or 2D-centric layouts provide insufficient geometric detail for handling challenges such as occlusion, object permanence, and geometry-aware planning. In contrast, volumetric or 3D occupancy representations such as neural fields [17] and structured voxel grids [18] provide geometric structure that better supports forecasting and control. Taken together, these points establish temporal modeling and spatial representation as core design dimensions that fundamentally influence the predictive horizon, physical fidelity, and downstream performance of embodied agents.
@@
Several recent surveys have organized the rapidly growing literature on world models. Overall, these surveys follow two main approaches. The first is a function-oriented perspective. For example, Ding et al. [4] categorized relevant works based on the two core functions of understanding and prediction, while Zhu et al. [19] presented a framework based on the core capabilities of world models. The second approach is application-driven, focusing on specific domains such as autonomous driving. Notably, Guan et al. [20] and Feng et al. [21] provided overviews of world model techniques for autonomous driving.
@@
Fig. 1 presents an overview of the structure and taxonomy of this paper. We begin in §2 by outlining the core concepts and theoretical foundations of world models. §3 introduces our three-axis taxonomy and maps representative methods onto this framework. §4 surveys the datasets and evaluation metrics used for training and assessment. §5 offers a quantitative comparison of state-of-the-art models. §6 discusses open challenges and promising research directions, and §7 concludes the survey.
§1 Introduction §2 Background §3 Taxonomy §4 Data and Metrics §5 Performance §6 Challenges and Trends §7 Conclusion Refer to captionCore Concepts 2.1)Refer to captionDecision-Coupled 3.1)Refer to captionGeneral-Purpose 3.2) Sequential 3.1.1 & §3.2.1) Refer to captionRecurrent StructureRefer to captionAutoregressive Methods Global 3.1.2 & §3.2.2) Refer to captionGlobal PredictionRefer to captionMasked JEPA Spatial Representation Refer to captionGlobal Latent VectorRefer to captionToken Feature SequenceRefer to captionSpatial Latent GridRefer to captionDecomposed Rendering Representation
Figure 1: Structure of this survey. The figure classifies world models along three axes and illustrates representative methods for each, providing a unified view of the field. Figure design inspired in part by [14, 12, 22, 23].
2 Background
2.1 Core Concepts
As discussed in §1, world models function as an internal simulator of environmental dynamics. Its functionality rests on three aspects:
@@ 2.1 Core Concepts
These three pillars provide the conceptual foundation for the taxonomy introduced in §3 and are formalized in the mathematical framework that follows.
2.2 Mathematical Formulation of World Models
We formalize the environment interaction as a POMDP [24]. For notational consistency, we define a null initial action $a_{0}$ at $t=0$, which allows the dynamics to be written uniformly. At each step $t\geq 1$, the agent receives an observation $o_{t}$ and takes an action $a_{t}$, while the true state $s_{t}$ remains unobserved. To handle this partial observability, world models infer a learned latent state $z_{t}$ using a one-step filtering posterior, where the previous latent state $z_{t-1}$ is assumed to summarize the relevant history. Finally, $z_{t}$ is used to reconstruct $o_{t}$:
@@ 2.2 Mathematical Formulation of World Models
$p_{\theta}(o_{1:T},z_{0:T}\!\mid\!a_{0:T-1})\!=\!p_{\theta}(z_{0})\!\prod_{t=1}^{T}\!p_{\theta}(z_{t}\!\mid\!z_{t-1},a_{t-1})p_{\theta}(o_{t}\!\mid\!z_{t}),$ (2)
@@
Modern world models thus adopt a reconstruction–regularization training paradigm: the likelihood term $\log p_{\theta}(o_{t}\!\mid\!z_{t})$ encourages faithful observation prediction, and KL regularization terms align the filtered posterior $q_{\phi}(z_{t}\!\mid\!z_{t-1},a_{t-1},o_{t})$ with the dynamics prior $p_{\theta}(z_{t}\!\mid\!z_{t-1},a_{t-1})$. Such world models can be instantiated with recurrent models [25, 26, 27], Transformer-based architectures [28, 29, 30], or diffusion-based decoders [31, 32, 33, 34, 35]. In all cases, the learned latent trajectory $z_{1:T}$ serves as a compact, predictive memory to support downstream policy optimization, model-predictive control, and counterfactual reasoning in embodied AI.
3 Taxonomy
@@
    Decomposed Rendering Representation involves decomposing 3D scenes into a set of learnable primitives, such as 3D Gaussian Splatting (3DGS) [36] or Neural Radiance Fields (NeRF) [37], and then using differentiable rendering to achieve high-fidelity novel view synthesis.
@@
PlaNet[38] ICML’19 Dec/Seq/GLV RSSM 1
Dreamer[10] ICLR’20 Dec/Seq/GLV RSSM 3
GLAMOR[39] ICLR’21 Dec/Seq/GLV IDM 2
DreamerV2[11] ICLR’21 Dec/Seq/GLV RSSM 2
TransDreamer[28] arXiv’22 Dec/Seq/GLV TSSM 4
Iso-Dream[40] NeurIPS’22 Dec/Seq/GLV IDM 4
MWM[41] CoRL’22 Dec/Seq/TFS RSSM 3
Inner Monologue[42] CoRL’22 Dec/Seq/TFS CoT 3
DayDreamer[43] CoRL’22 Dec/Seq/GLV RSSM 4
TWM[29] ICLR’23 Dec/Seq/TFS Transformer 1
IRIS[44] ICLR’23 Dec/Seq/TFS Transformer 1
WorldDreamer[45] arXiv’24 Gen/Glo/TFS Transformer 4
Statler[46] ICRA’24 Dec/Seq/TFS LLM 2
Pandora[47] arXiv’24 Gen/Seq/TFS Video Diffusion 2
DWL[48] RSS’24 Dec/Seq/GLV MLP 4
RoboDreamer[49] ICML’24 Dec/Glo/TFS IDM 2
Genie[50] ICML’24 Gen/Seq/TFS Transformer 3
V-JEPA[51] TMLR’24 Gen/Glo/TFS JEPA 6
PreLAR[52] ECCV’24 Dec/Seq/GLV RSSM 3
ManiGaussian[53] ECCV’24 Dec/Seq/DRR 3DGS 1
ECoT[54] CoRL’24 Dec/Glo/TFS CoT 3
VidMan[55] NeurIPS’24 Dec/Glo/TFS IDM 4
iVideoGPT[56] NeurIPS’24 Gen/Seq/TFS Transformer 6
EnerVerse[34] arXiv’25 Dec/Seq/SLG Video Diffusion 4
GLAM[57] AAAI’25 Dec/Seq/GLV Mamba 1
NavCoT[58] TPAMI’25 Dec/Seq/TFS CoT 4
DreamerV3[12] Nature’25 Dec/Seq/GLV RSSM 8
MineWorld[59] arXiv’25 Dec/Seq/TFS Transformer 1
DreMa[60] ICLR’25 Dec/Seq/DRR 3DGS 2
S2-SSM[61] arXiv’25 Gen/Seq/TFS Mamba 1
RLVR-World[62] arXiv’25 Gen/Seq/TFS RLVR 3
StateSpaceDiffuser[63] arXiv’25 Gen/Seq/TFS Mamba 2
DeepVerse[64] arXiv’25 Gen/Seq/TFS DiT 1
ORV[65] arXiv’25 Gen/Glo/SLG DiT 4
V-JEPA 2[14] arXiv’25 Gen/Glo/TFS JEPA 15
NWM[66] CVPR’25 Dec/Seq/TFS DiT 6
WorldVLA[67] arXiv’25 Dec/Seq/TFS Transformer 1
World4Omni[68] arXiv’25 Gen/Seq/TFS VLM 2
Dyn-O[69] arXiv’25 Dec/Seq/TFS Mamba 1
DINO-WM[70] ICML’25 Dec/Seq/SLG Transformer 3
EVA[71] ICML’25 Gen/Seq/TFS RoG 4
AdaWorld[72] ICML’25 Gen/Seq/TFS Video Diffusion 6
MindJourney[73] arXiv’25 Gen/Seq/SLG VLM 2
GAF[74] arXiv’25 Dec/Seq/DRR 4DGS 1
Yume[75] arXiv’25 Gen/Seq/TFS DiT 1
villa-X[76] arXiv’25 Dec/Glo/TFS IDM 5
AETHER[77] ICCV’25 Gen/Glo/SLG DiT 6
TesserAct[78] ICCV’25 Dec/Glo/SLG IDM 4
MineDreamer[79] IROS’25 Dec/Seq/TFS CoI 3
ManiGaussian++[80] IROS’25 Dec/Seq/DRR 3DGS 2
  • 1 Taxonomy: Abbreviations for the taxonomy categories defined in §3.
@@
MILE[81] NeurIPS’22 Dec/Seq/GLV RSSM 1
Copilot4D[82] ICLR’24 Gen/Seq/SLG Video Diffusion 3
SEM2[83] TITS’24 Dec/Seq/GLV RSSM 1
MagicDrive3D[84] arXiv’24 Gen/Glo/DRR 3DGS 1
OccSora[85] arXiv’24 Gen/Glo/SLG Diffusion 2
Delphi[86] arXiv’24 Gen/Seq/SLG Video Diffusion 1
DriveWorld[87] CVPR’24 Dec/Seq/SLG RSSM 2
Drive-WM[88] CVPR’24 Dec/Glo/SLG Video Diffusion 1
ViDAR[89] CVPR’24 Gen/Seq/SLG Transformer 1
GenAD[90] CVPR’24 Gen/Seq/TFS Video Diffusion 4
OccLLaMA[18] arXiv’24 Dec/Seq/SLG Transformer 3
DriveDreamer[91] ECCV’24 Dec/Seq/SLG GRU 1
GenAD[92] ECCV’24 Dec/Seq/SLG GRU 1
OccWorld[93] ECCV’24 Dec/Seq/SLG Transformer 2
DOME[94] arXiv’24 Gen/Seq/SLG DiT 2
TOKEN[95] CoRL’24 Dec/Glo/TFS Transformer 2
Vista[96] NeurIPS’24 Gen/Seq/SLG Video Diffusion 4
DriveDreamer-2[97] AAAI’25 Gen/Glo/SLG Video Diffusion 1
DTT[98] arXiv’25 Dec/Seq/DRR Transformer 2
DynamicCity[99] ICLR’25 Gen/Glo/SLG DiT 4
LidarDM[100] ICRA’25 Gen/Seq/SLG Diffusion 3
FutureSightDrive[101] arXiv’25 Dec/Seq/TFS CoT(VLM) 3
GEM[102] CVPR’25 Gen/Seq/SLG Video Diffusion 1
GaussianWorld[103] CVPR’25 Gen/Seq/DRR Transformer 2
MaskGWM[104] CVPR’25 Gen/Glo/TFS DiT 3
DriveDreamer4D[105] CVPR’25 Gen/Glo/DRR 4DGS 3
ReconDreamer[106] CVPR’25 Gen/Glo/DRR 3DGS 3
WoTE[107] ICCV’25 Dec/Seq/SLG Transformer 2
HERMES[108] ICCV’25 Gen/Glo/SLG LLM 4
InfiniCube[22] ICCV’25 Gen/Seq/DRR 3DGS 1
DriVerse[109] ACMMM’25 Gen/Seq/TFS DiT 2
  • 1 Taxonomy: Abbreviations for the taxonomy categories defined in §3.
@@
3.1 Decision-Coupled World Models
3.1.1 Sequential Simulation and Inference

### v2

@@ I Introduction
Embodied AI aims to equip agents to perceive complex, multimodal environments, act within them, and anticipate how their actions will alter future world states [batra2020rearrangement, gupta2024essential]. Central to this capability is the world model, an internal simulator that captures environment dynamics to support both forward and counterfactual rollouts for perception, prediction, and decision making [liu2025aligning, ding2024understanding]. This survey focuses on world models that yield actionable predictions for embodied agents, distinguishing them from static scene descriptors or purely generative visual models that do not capture controllable dynamics.
@@
Cognitive science suggests humans construct internal models of the world by integrating sensory inputs. These models do not merely predict and simulate future events but also shape perception and guide action [clark1998being, barsalou1999perceptions, friston2010free]. Motivated by this view, early AI research on world models was rooted in model-based reinforcement learning (RL), where latent state-transition models were used to improve sample efficiency and planning performance [fung2025embodied]. The seminal work of Ha and Schmidhuber [ha2018recurrent] crystallized the term world model and inspired the Dreamer series [hafner2020dream, hafner2021mastering, hafner2025mastering], highlighting how learned dynamics can drive imagination-based policy optimization. More recently, advances in large-scale generative modeling and multimodal learning have expanded world models beyond their initial focus on policy learning into general-purpose environment simulators capable of high-fidelity future prediction, exemplified by models like Sora [brooks2024video] and V-JEPA 2 [assran2025v]. This expansion has diversified functional roles, temporal modeling strategies, and spatial representations, while introducing inconsistencies in terminology and taxonomy across sub-communities.
@@
Faithfully capturing environment dynamics requires addressing both the temporal evolution of states and the spatial encoding of scenes [liu2025aligning]. Long-horizon rollouts are susceptible to error accumulation, which establishes coherence as a central challenge in video prediction and policy imagination [venkatraman2015improving, asadi2018lipschitz]. Similarly, coarse or 2D-centric layouts provide insufficient geometric detail for handling challenges such as occlusion, object permanence, and geometry-aware planning. In contrast, volumetric or 3D occupancy representations such as neural fields [agro2024uno] and structured voxel grids [wei2024occllama] provide geometric structure that better supports forecasting and control. Taken together, these points establish temporal modeling and spatial representation as core design dimensions that fundamentally influence the predictive horizon, physical fidelity, and downstream performance of embodied agents.
@@
Several recent surveys have organized the rapidly growing literature on world models. Overall, these surveys follow two main approaches. The first is a function-oriented perspective. For example, Ding et al. [ding2024understanding] categorized relevant works based on the two core functions of understanding and prediction, while Zhu et al. [zhu2024sora] presented a framework based on the core capabilities of world models. The second approach is application-driven, focusing on specific domains such as autonomous driving. Notably, Guan et al. [guan2024world] and Feng et al. [feng2025survey] provided overviews of world model techniques for autonomous driving.
@@
Fig. 1 presents an overview of the structure and taxonomy of this paper. We begin in §II by outlining the core concepts and theoretical foundations of world models. §III introduces our three-axis taxonomy and maps representative methods onto this framework. §IV surveys the datasets and evaluation metrics used for training and assessment. §V offers a quantitative comparison of state-of-the-art models. §VI discusses open challenges and promising research directions, and §VII concludes the survey.
§I Introduction §II Background §III Taxonomy §IV Data and Metrics §V Performance §VI Challenges and Trends §VII Conclusion Refer to captionCore Concepts II-A)Refer to captionDecision-Coupled III-A)Refer to captionGeneral-Purpose III-B) Sequential III-A1 & §III-B1) Refer to captionRecurrent StructureRefer to captionAutoregressive Methods Global III-A2 & §III-B2) Refer to captionGlobal PredictionRefer to captionMasked JEPA Spatial Representation Refer to captionGlobal Latent VectorRefer to captionToken Feature SequenceRefer to captionSpatial Latent GridRefer to captionDecomposed Rendering Representation
Figure 1: Structure of this survey. The figure classifies world models along three axes and illustrates representative methods for each, providing a unified view of the field. Figure design inspired in part by [assran2025v, hafner2025mastering, lu2025infinicube, chen2024drivinggpt].
II Background
II-A Core Concepts
As discussed in §I, a world model functions as an internal simulator of environmental dynamics. Its functionality rests on three aspects:
@@ II-A Core Concepts
These three pillars provide the conceptual foundation for the taxonomy introduced in §III and are formalized in the mathematical framework that follows.
II-B Mathematical Formulation of World Models
We formalize the environment interaction as a POMDP [smallwood1973optimal]. For notational consistency, we define a null initial action $a_{0}$ at $t=0$, which allows the dynamics to be written uniformly. At each step $t\geq 1$, the agent receives an observation $o_{t}$ and takes an action $a_{t}$, while the true state $s_{t}$ remains unobserved. To handle this partial observability, world models infer a learned latent state $z_{t}$ using a one-step filtering posterior, where the previous latent state $z_{t-1}$ is assumed to summarize the relevant history. Finally, $z_{t}$ is used to reconstruct $o_{t}$:
@@ II-B Mathematical Formulation of World Models
$p_{\theta}(o_{1:T},z_{0:T}\!\mid\!a_{0:T-1})\!=\!p_{\theta}(z_{0})\!\prod_{t=1}^{T}\!p_{\theta}(z_{t}\!\mid\!z_{t-1},a_{t-1})p_{\theta}(o_{t}\!\mid\!z_{t}).$ (2)
@@
Modern world models thus adopt a reconstruction–regularization training paradigm: the likelihood term $\log p_{\theta}(o_{t}\!\mid\!z_{t})$ encourages faithful observation prediction, and KL regularization terms align the filtered posterior $q_{\phi}(z_{t}\!\mid\!z_{t-1},a_{t-1},o_{t})$ with the dynamics prior $p_{\theta}(z_{t}\!\mid\!z_{t-1},a_{t-1})$. Such world models can be instantiated with recurrent models [sun2024learning, zhai2025recurrent, aljalbout2025accelerating], Transformer-based architectures [chen2022transdreamer, robine2023transformerbased, zhang2023storm], or diffusion-based decoders [zhu2025unified, qi2025strengthening, guo2025flowdreamer, huang2025enerverse, alhaija2025cosmos]. In all cases, the learned latent trajectory $z_{1:T}$ serves as a compact, predictive memory to support downstream policy optimization, model-predictive control, and counterfactual reasoning in embodied AI.
III Taxonomy
@@
    Decomposed Rendering Representation involves decomposing 3D scenes into a set of learnable primitives, such as 3D Gaussian Splatting (3DGS) [kerbl20233d] or Neural Radiance Fields (NeRF) [mildenhall2020nerf], and then using differentiable rendering to achieve high-fidelity novel view synthesis.
@@
PlaNet[hafner2019learning] ICML’19 Dec/Seq/GLV RSSM 1
Dreamer[hafner2020dream] ICLR’20 Dec/Seq/GLV RSSM 3
GLAMOR[paster2021planning] ICLR’21 Dec/Seq/GLV IDM 2
DreamerV2[hafner2021mastering] ICLR’21 Dec/Seq/GLV RSSM 2
TransDreamer[chen2022transdreamer] arXiv’22 Dec/Seq/GLV TSSM 4
Iso-Dream[pan2022iso] NeurIPS’22 Dec/Seq/GLV IDM 4
MWM[seo2023masked] CoRL’22 Dec/Seq/TFS RSSM 3
Inner Monologue[huang2023inner] CoRL’22 Dec/Seq/TFS CoT 3
DayDreamer[wu2023daydreamer] CoRL’22 Dec/Seq/GLV RSSM 4
TWM[robine2023transformerbased] ICLR’23 Dec/Seq/TFS Transformer 1
IRIS[micheli2023transformers] ICLR’23 Dec/Seq/TFS Transformer 1
WorldDreamer[wang2024worlddreamer] arXiv’24 Gen/Glo/TFS Transformer 4
Statler[yoneda2024statler] ICRA’24 Dec/Seq/TFS LLM 2
Pandora[xiang2024pandora] arXiv’24 Gen/Seq/TFS Video Diffusion 2
DWL[gu2024advancing] RSS’24 Dec/Seq/GLV MLP 4
RoboDreamer[zhou2024robodreamer] ICML’24 Dec/Glo/TFS IDM 2
Genie[bruce2024genie] ICML’24 Gen/Seq/TFS Transformer 3
V-JEPA[bardes2024revisiting] TMLR’24 Gen/Glo/TFS JEPA 6
PreLAR[zhang2024prelar] ECCV’24 Dec/Seq/GLV RSSM 3
ManiGaussian[lu2024manigaussian] ECCV’24 Dec/Seq/DRR 3DGS 1
ECoT[zawalski2025robotic] CoRL’24 Dec/Glo/TFS CoT 3
VidMan[wen2024vidman] NeurIPS’24 Dec/Glo/TFS IDM 4
iVideoGPT[wu2024ivideogpt] NeurIPS’24 Gen/Seq/TFS Transformer 6
EnerVerse[huang2025enerverse] arXiv’25 Dec/Seq/SLG Video Diffusion 4
GLAM[he2025glam] AAAI’25 Dec/Seq/GLV Mamba 1
NavCoT[lin2025navcot] TPAMI’25 Dec/Seq/TFS CoT 4
DreamerV3[hafner2025mastering] Nature’25 Dec/Seq/GLV RSSM 8
MineWorld[guo2025mineworld] arXiv’25 Dec/Seq/TFS Transformer 1
DreMa[barcellona2025dream] ICLR’25 Dec/Seq/DRR 3DGS 2
S2-SSM[petri2025learning] arXiv’25 Gen/Seq/TFS Mamba 1
RLVR-World[wu2025rlvr] arXiv’25 Gen/Seq/TFS RLVR 3
StateSpaceDiffuser[savov2025statespacediffuser] arXiv’25 Gen/Seq/TFS Mamba 2
DeepVerse[chen2025deepverse] arXiv’25 Gen/Seq/TFS DiT 1
ORV[yang2025orv] arXiv’25 Gen/Glo/SLG DiT 4
V-JEPA 2[assran2025v] arXiv’25 Gen/Glo/TFS JEPA 15
NWM[bar2025navigation] CVPR’25 Dec/Seq/TFS DiT 6
WorldVLA[cen2025worldvla] arXiv’25 Dec/Seq/TFS Transformer 1
World4Omni[chen2025world4omni] arXiv’25 Gen/Seq/TFS VLM 2
Dyn-O[wang2025dyn] arXiv’25 Dec/Seq/TFS Mamba 1
DINO-WM[zhou2025dino] ICML’25 Dec/Seq/SLG Transformer 3
EVA[chi2025empowering] ICML’25 Gen/Seq/TFS RoG 4
AdaWorld[gao2025adaworld] ICML’25 Gen/Seq/TFS Video Diffusion 6
MindJourney[yang2025mindjourney] arXiv’25 Gen/Seq/SLG VLM 2
GAF[chai2025gaf] arXiv’25 Dec/Seq/DRR 4DGS 1
Yume[mao2025yume] arXiv’25 Gen/Seq/TFS DiT 1
villa-X[chen2025villa] arXiv’25 Dec/Glo/TFS IDM 5
AETHER[team2025aether] ICCV’25 Gen/Glo/SLG DiT 6
TesserAct[zhen2025tesseract] ICCV’25 Dec/Glo/SLG IDM 4
MineDreamer[zhou2025minedreamer] IROS’25 Dec/Seq/TFS CoI 3
ManiGaussian++[yu2025manigaussian++] IROS’25 Dec/Seq/DRR 3DGS 2
  • 1 Taxonomy: Abbreviations for the taxonomy categories defined in §III.
@@
MILE[hu2022model] NeurIPS’22 Dec/Seq/GLV RSSM 1
Copilot4D[zhang2024copilot4d] ICLR’24 Gen/Seq/SLG Video Diffusion 3
SEM2[gao2024enhance] TITS’24 Dec/Seq/GLV RSSM 1
MagicDrive3D[gao2024magicdrive3d] arXiv’24 Gen/Glo/DRR 3DGS 1
OccSora[wang2024occsora] arXiv’24 Gen/Glo/SLG Diffusion 2
Delphi[ma2024unleashing] arXiv’24 Gen/Seq/SLG Video Diffusion 1
DriveWorld[min2024driveworld] CVPR’24 Dec/Seq/SLG RSSM 2
Drive-WM[wang2024driving] CVPR’24 Dec/Glo/SLG Video Diffusion 1
ViDAR[Yang2024visual] CVPR’24 Gen/Seq/SLG Transformer 1
GenAD[yang2024generalized] CVPR’24 Gen/Seq/TFS Video Diffusion 4
OccLLaMA[wei2024occllama] arXiv’24 Dec/Seq/SLG Transformer 3
DriveDreamer[wang2024drivedreamer] ECCV’24 Dec/Seq/SLG GRU 1
GenAD[zheng2024genad] ECCV’24 Dec/Seq/SLG GRU 1
OccWorld[zheng2024occworld] ECCV’24 Dec/Seq/SLG Transformer 2
DOME[gu2024dome] arXiv’24 Gen/Seq/SLG DiT 2
TOKEN[tian2025tokenize] CoRL’24 Dec/Glo/TFS Transformer 2
Vista[gao2024vista] NeurIPS’24 Gen/Seq/SLG Video Diffusion 4
DriveDreamer-2[zhao2025drivedreamer] AAAI’25 Gen/Glo/SLG Video Diffusion 1
DTT[xu2025delta] arXiv’25 Dec/Seq/DRR Transformer 2
DynamicCity[bian2025dynamiccity] ICLR’25 Gen/Glo/SLG DiT 4
LidarDM[zyrianov2025lidardm] ICRA’25 Gen/Seq/SLG Diffusion 3
FutureSightDrive[zeng2025futuresightdrive] arXiv’25 Dec/Seq/TFS CoT(VLM) 3
GEM[hassan2025gem] CVPR’25 Gen/Seq/SLG Video Diffusion 1
GaussianWorld[zuo2025gaussianworld] CVPR’25 Gen/Seq/DRR Transformer 2
MaskGWM[ni2025maskgwm] CVPR’25 Gen/Glo/TFS DiT 3
DriveDreamer4D[zhao2025drivedreamer4d] CVPR’25 Gen/Glo/DRR 4DGS 3
ReconDreamer[ni2025recondreamer] CVPR’25 Gen/Glo/DRR 3DGS 3
WoTE[li2025end] ICCV’25 Dec/Seq/SLG Transformer 2
HERMES[zhou2025hermes] ICCV’25 Gen/Glo/SLG LLM 4
InfiniCube[lu2025infinicube] ICCV’25 Gen/Seq/DRR 3DGS 1
DriVerse[li2025driverse] ACMMM’25 Gen/Seq/TFS DiT 2
  • 1 Taxonomy: Abbreviations for the taxonomy categories defined in §III.
@@
III-A Decision-Coupled World Models
III-A1 Sequential Simulation and Inference